Recent Log Entries

Communications error

Log Number: 3738831Logbook(s): Entry Makers: lawBody: 

Shunt MXR2S01 Intermittant mismatch and comm error. Will not clear. Disabled in alarm handler until corrected.

Arc 2 Faulted Off

Log Number: 3738830Logbook(s): Entry Makers: lawBody: 

Arc 2 tripped off. Fault screens below. Fault cycles through "No response" and "bad message length".

Operator SIGN IN

Log Number: 3738829Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Surles-Law
DATE:             11/03/2019
SHIFT:               Owl

The operator has read the following:

Follow-up Re: NL low heater margin vs. cold compressor flow

Log Number: 3738828Logbook(s): Body: 

LEM at 1830 raised North Linac C1O0's and a few others up that created the heat load. Ops turned the C100's back to where they are running since Saturday and not doing any more gradient changes until Monday. We will monitor next weeks changes and suspect we will be trying to get the last of the C100 running and at higher gradients.

1L08-3 Heating and Zone Drop

Log Number: 3738823Logbook(s): Entry Makers: kernsBody: 

Was monitoring this zone since turn on last week and didn't lower to 10Mv/m before LEM today raised it to 11Mv/m and heating was taking the zone down. Good work C. Humphrey and thanks. Set to 10Mv/m in ODVH or drop gradient lower if needed.

CREW CHIEF SIGN IN

Log Number: 3738821Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 11-02-19
Time : 23:05:53
Day : Saturday
Shift : Owl

Crew Chief : Frierson
Program Deputy: Michaud
Team : X
Arm : RADCON
Designated SSO: Surles-Law
LERF Operator:
Operators : Surles-Law

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

1L08 Cav 3 Heat Issue

Log Number: 3738817Logbook(s): Entry Makers: deirBody: 

Cavity three on 1L08 seems to have a heat issue. After all eight cavities had been turned on I proceeded to walk up cavity three to its ODVH; however, GSET increments of 0.100 made relatively drastic changes in the liquid level, which continued to rise resulting in a trip of the whole module.

RF Checks 11/02/19

Log Number: 3738816Logbook(s): Body: 

1L20 cavities 1 and 2 brought back into RF after LEM left them off. They were still marked bad and being
monitored after repairs during commissioning of new cryo-module. Marked as good since they are tuned and look
good since Friday.
1L09-5 Was being monitored for CWWT at 8Mv/m and IR voltage seems stable at ~1.6v. This one was also marked
good in cavity history so it stays on to be monitored.
Checked up on some repairs Friday and they all stayed up after LEM was performed at 1830. Looking good and hope the best for the rest !!!

1L21-8 KMAI

Log Number: 3738815Logbook(s): Entry Makers: kernsBody: 

KMAI holding steady at ~1.72ma with 1.2kv since Fridays repairs.

MAC Access Issues

Log Number: 3738810Logbook(s): Entry Makers: deirBody: 

I don't have channel access to ioc's related to 0L04 and am also not permitted to Set OFF the RF Auto Reset Panel for individual cryo modules.

CAV History Problem

Log Number: 3738809Logbook(s): Entry Makers: chumphryBody: 

I am not able to pull up CAV History on any zone from CAV History button

NL low heater margin vs. cold compressor flow

Log Number: 3738807Logbook(s): Body: 

At 1900, the Cryo Coordinator reported low NL heater margin, which was - 10 watts per cryomodule.
I remotely logged in to see recent increase in RF power and increase in NL helium vapor pressure to 0.0397 atm.
To reduce NL helium vapor pressure, I increased SCN cold compressor flow from 195 g/s to 199 g/s, which would eventually restore heater margin.

Pages